Calculate the energy of a 1 MeV gamma-ray photon after Compton scattering through 90 degrees


1
Expert's answer
2021-06-14T14:10:11-0400

Gives


Energy of gamma photon

"E_\gamma=1MeV=1.6\times10^{-13}"

"\phi=90°"

"E_e=0.511MeV"

"K=\frac{hv\alpha(1-cos\phi)}{1+\alpha(1-cos\phi)}"

Where

"\alpha=\frac{hv}{m_0c^2}=\frac{1MeV}{0.511MeV}=1.96"

put alpha value


"K=\frac{1.6\times10^{-13}\times1.96\times(1-cos90°)}{1+1.96\times(1-cos90°)}=1.05\times10^{-13}J"